# CodeAgent: Enhancing Code Generation with Tool-Integrated Agent Systems for Real-World Repo-level Coding Challenges

### v1

@@ Abstract
Large Language Models (LLMs) have shown promise in automated code generation but typically excel only in simpler tasks such as generating standalone code units. Real-world software development, however, often involves complex code repositories (named repo) with complex dependencies and extensive documentation. To fill this gap, our research pivots towards evaluating LLMs in a more realistic setting – real-world repo-level code generation. We introduce CodeAgentBench, a manually curated benchmark for repo-level code generation. This benchmark comprises five high-quality Python projects, encompassing a total of 101 samples. We assess nine leading LLMs on repo-level tasks and observe a decline in their performance. To tackle this, we present CodeAgent, a novel LLM-based agent framework that employs external tools for effective repo-level code generation. CodeAgent integrates five programming tools, enabling interaction with software artifacts for information retrieval, code symbol navigation, and code testing. We implement four agent strategies to optimize these tools’ usage. Our experiments on CodeAgentBench show that CodeAgent enhances LLM performance significantly, with improvements ranging from 18.1% to 250%. Further tests on the HumanEval benchmark confirm CodeAgent’s adaptability and efficacy across various code generation tasks. Notably, CodeAgent outperforms commercial products like Github Copilot, showcasing superior accuracy and efficiency. These results demonstrate CodeAgent’s robust capabilities in code generation, highlighting its potential for real-world repo-level coding challenges.  [^1] Code, data, and more case studies are open-sourced on Github: https://github.com/zkcpku/CodeAgent
@@ 1 Introduction
Code generation automatically generates programs for the natural language (NL) requirement. Recent years have seen a trend in tackling code generation tasks with large language models (LLMs), such as CodeLlama Rozière et al. (2023), StarCoder Li et al. (2023), and DeepSeekCoder DeepSeek (2023). To evaluate the effectiveness of LLMs on code generation, many efforts have been performed Zhang et al. (2023b); Luo et al. (2023); Zheng et al. (2023). They demonstrate LLMs have impressive code generation abilities, e.g., CodeLlama Rozière et al. (2023) achieves 55% pass rate on the MBPP benchmark Austin et al. (2021). Despite achieving satisfactory performances, they mainly focus on simple generation scenarios including statement-level and function-level code generation. Statement-level code generation Iyer et al. (2018); Athiwaratkun et al. (2022) aims to output statement-specific source codes. Function-level code generation Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021) predicts independent code that only invokes built-in functions and APIs from third-party libraries. For both scenarios, on the one hand, the length of the generated code is rather short. For instance, the average line of each target code is 11.5 on the most popular benchmark HumanEval Chen et al. (2021). On the other hand, they only generate standalone code units. However, more than 70% functions in the open-source projects are non-standalone Yu et al. (2023). Developers typically write programs based on specific code environments that usually have intricate contextual dependencies.
To address the gap in understanding LLMs’ performance in practical scenarios, we attempt to investigate LLMs in a more pragmatic scenario, i.e., repo-level code generation. We manually construct CodeAgentBench, a benchmark specifically for repo-level code generation, comprising five diverse Python code repositories with a total of 101 functions sourced from Github  [^2] https://github.com/. It provides rich information about the repository, such as documentation and contextual dependency, to help LLMs better understand it. Besides, each repository has a self-contained test suite to efficiently verify the correctness of generated codes. Given a requirement, LLMs generate desired programs that not only meet the requirement but also ensure seamless integration with the current project repository. When tested with CodeAgentBench, we observe that even the most powerful model, GPT-4 GPT-4 (2023), only achieves 21.8% pass rate, indicating a decline in performance compared to standalone code generation. These models struggle with handling complex code tasks, often leading to inaccurate code outputs. Further details and examples are provided in Section 6.2 and Figure 5.
@@
To alleviate this problem, we follow the research line of LLM-based agents Palo et al. (2023); Wang et al. (2023a); Xi et al. (2023) and introduce a novel LLM-based agent framework CodeAgent that leverages external tools to help LLMs in repo-level code generation. With five programming tools, CodeAgent is capable of interacting with the software artifacts, including retrieving useful information, finding existing code symbols in the repository, and handling essential code testing. To guide LLMs to efficiently use tools, we draw on four agent strategies including ReAct, Tool-Planning, OpenAI function, and Rule-based form. Based on agent strategies, LLMs can automatically select suitable tools for each repo-level task, finally provide a comprehensive response. To our knowledge, we are the first to adopt and optimize a agent-based method for complex repo-level code generation tasks.
@@
We conduct extensive experiments for evaluation. We apply CodeAgent to nine powerful open-source and closed-source LLMs with parameter sizes ranging from 13B to 175B to show the universality. Compared to directly generating from LLMs, experimental results on CodeAgentBench reveal that CodeAgent achieves performance improvements ranging from 18.1% to an extraordinary 250% across various LLMs. Further evaluations on well-known function-level benchmark HumanEval Chen et al. (2021) confirm CodeAgent’s versatility and effectiveness in diverse code generation tasks. Remarkably, when compared with commercial products like Github Copilot, CodeAgent stands out, demonstrating superior accuracy. These findings highlight the robust practical capabilities of CodeAgent in code generation community, underscoring its potential to evolve real-world repo-level coding challenges.
In summary, we make the following main contributions:
@@
  • We make an attempt to investigate repo-level code generation, which has crucial worth for understanding LLMs’ performance in practical scenarios.
  • We construct CodeAgentBench, a repo-level code generation benchmark. It has high-quality code repositories and diverse topics.
@@
  • We propose CodeAgent that leverages tools to tackle repo-level code generation. We explore five tools to help LLMs complete the whole generation process and draw on four agent strategies to automatically select suitable tools and generation formats.
@@
  • Experimental results on nine LLMs demonstrate our CodeAgent’s versatility and effectiveness in diverse code generation tasks, highlighting its potential for real-world repository-level coding challenges.

### v2

@@ Abstract
Large Language Models (LLMs) have shown promise in automated code generation but typically excel only in simpler tasks such as generating standalone code units. However, real-world software development often involves complex code repositories with complex dependencies and extensive documentation. To enable LLMs to handle these real-world repo-level code generation, we present CodeAgent, a novel LLM-based agent framework that employs external tools for effective repo-level code generation. CodeAgent integrates five programming tools, enabling interaction with software artifacts for information retrieval, code implementation, and code testing. We implement four agent strategies to optimize these tools’ usage. To the best of our knowledge, CodeAgent is the first agent framework specifically for repo-level code generation. In order to measure the effectiveness of our method at the repository level, we design a repo-level benchmark CodeAgentBench. The performance on this benchmark shows a significant improvement brought by our method, with improvements in pass rate ranging from 2.0 to 15.8. Further tests on the HumanEval benchmark confirm CodeAgent’s adaptability and efficacy across various code generation tasks. Notably, CodeAgent outperforms commercial products like GitHub Copilot, showcasing superior accuracy and efficiency. These results demonstrate CodeAgent’s robust capabilities in code generation, highlighting its potential for real-world repo-level coding challenges.
@@ 1 Introduction
Code generation automatically generates programs for the natural language (NL) requirement. Recent years have seen a trend in tackling code generation tasks with large language models (LLMs), such as Code Llama Rozière et al. (2023), StarCoder Li et al. (2023), and DeepSeekCoder DeepSeek (2023). Many efforts have been performed Zhang et al. (2023b); Luo et al. (2023); Zheng et al. (2023) and shown impressive code generation abilities.
@@
Despite achieving satisfactory performances, these studies mainly focus on simple generation scenarios including statement-level and function-level code generation. Statement-level code generation Iyer et al. (2018); Athiwaratkun et al. (2022) aims to output statement-specific source codes. Function-level code generation Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021) predicts independent code that only invokes built-in functions and APIs from third-party libraries. For both scenarios, the length of the generated code is rather short, and they only generate standalone code units. However, more than 70% functions in the open-source projects are non-standalone Yu et al. (2023). Developers typically write programs based on specific code environments, generally referring to code repositories. These repo-level code snippets usually have intricate contextual dependencies, which is too complex for existing LLMs to handle and generate Li et al. (2024).
@@
To enhance the efficacy of LLMs in repo-level code generation tasks, we draw inspiration from human programming practices. Developers typically employ a variety of tools to aid in complex programming. For instance, they might utilize search engines to explore key concepts or static analysis tools to identify pre-existing functions or classes. These tools are instrumental in the development of code projects. Embracing this idea, we propose a novel LLM-based agent framework CodeAgent that leverages external tools to help LLMs in repo-level code generation. With five programming tools, CodeAgent is capable of interacting with the software artifacts, including retrieving useful information, finding existing code symbols in the repository, and handling essential code testing. To guide LLMs to efficiently use tools, we draw on four agent strategies covering ReAct, Tool-Planning, OpenAIFunc, and Rule-based form. Based on agent strategies, LLMs can automatically select suitable tools for each repo-level task, finally providing a comprehensive response.
@@
In order to measure the effectiveness of our method at the code repository, we manually construct CodeAgentBench, a benchmark specifically for repo-level code generation with a total of 101 functions and classes sourced from real code projects. It provides rich information about the repository, such as documentation and contextual dependency, to help LLMs better understand it. We further conduct extensive experiments for evaluation. We apply CodeAgent to nine powerful open-source and closed-source LLMs with parameter sizes ranging from 13B to 175B to show the universality. Compared to directly generating from LLMs, experimental results on CodeAgentBench reveal that CodeAgent achieves significant improvements ranging from 2.0 to an extraordinary 15.8 across various LLMs. Further evaluations on well-known function-level benchmark HumanEval Chen et al. (2021) confirm CodeAgent’s versatility in diverse code generation tasks. Remarkably, when compared to commercial products like GitHub Copilot Dakhel et al. (2023), CodeAgent stands out, demonstrating superior accuracy. These findings highlight the robust practical capabilities of CodeAgent in the code generation community, underscoring its potential to evolve real-world repo-level coding challenges. We summarize our main contributions:
@@
  • We make an attempt to investigate repo-level code generation, which has crucial worth for understanding LLMs’ performance in practical code generation scenarios.
@@
  • We propose CodeAgent, an LLM-based agent framework for repo-level code generation. It develops five external programming tools to help LLMs complete the whole generation process and draw on four agent strategies to automatically optimize tools’ usage.
  • We construct CodeAgentBench, a repo-level code generation benchmark, which has high-quality code repositories and covers diverse topics.
@@
  • Experimental results on nine LLMs show CodeAgent’s versatility and effectiveness in diverse code generation tasks, highlighting its potential for resolving real-world repo-level coding challenges.